# DEPTH: Discourse Education through Pre-Training Hierarchically

### v1

@@ 3.1 Fine-Tuning Experiments
We follow up our pre-training experiments with Natural Language Understanding (NLU) downstream tasks. We evaluate our models on natural language inference (MNLI, [62]), sentiment analysis (SST2, [50]), and grammar (CoLA, [59]) within the GLUE benchmark [58]. We also use the DiscoEval suite [5] to evaluate models on their understanding of discourse. We use two tasks from this benchmark: Sentence Permutation (SP) and Discourse Coherence (DC). SP involves shuffling sentences, and predicting which of the shuffled sentences originally appeared first, while DC involves predicting whether or not the input paragraph was coherent. Finally, we measure our model’s generative abilities on the Natural Instructions (NI) dataset [37], which measures the ability of LMs to follow instructions, and served as a benchmark for NanoT5 [39].
Our experimental framework is inspired by Pythia [2], as we evaluate the performance of our models on downstream tasks from intermediate checkpoints. Unlike Pythia, we run evaluation with checkpoints from both T5 and DEPTH models, gathered at steps $\{2K,4K,\ldots,512K,1M\}$ in order to examine these models’ emergent capabilities. The exponential distance between these checkpoints allows us to scale intermediate checkpoint evaluation to much longer training runs  [^5] Evaluating intermediate checkpoint performance every 10,000 steps (as was done in Pythia) on datasets as large as MNLI is unfeasible with our limited computational resources..
@@ 4.1 C4 Pre-Training
During pre-training, we find that DEPTH consistently achieves a lower validation loss than a comparably trained T5 model. This is true for both FS pre-training, and CPT pre-training. Furthermore, when we isolate the reconstruction loss (the same objective used by T5, not including sentence tokens), we find that DEPTH outperforms T5 despite balancing an additional pre-training objective (Figure 3 for FS and Figure 3 for CPT). These results are consistent with the findings in SLM, where their model converged faster, and on fewer tokens than models such as BERT and T5.
@@
While we were not able to match the performance of the baseline model of Raffel et al. [45] (see Appendix A for speculations on why), we have obtained the lowest loss scores among PyTorch implementations of T5 models. Specifically, in our FS setting, we find that our randomly initialized T5 model outperforms the validation loss of NanoT5, achieving $1.65$ vs. $1.95$ at step 64,000.
@@ 4.2 GLUE Fine-Tuning
We found that over the course of FS pre-training, both models improved on GLUE tasks. However, T5’s improvement pattern was sporadic and slow when compared to that of DEPTH (top row of Figure 4). We found it difficult to replicate the results of the original T5 (both on GLUE tasks and the pre-training loss) as discussed in Appendix A. We project that with more substantial training (i.e., 1-3 million pre-training steps, and a larger batch size of $\geq 2048$ examples per batch, as in T5 [45]), DEPTH could match or exceed the performance of T5 and SLM on downstream tasks.
@@
Figure 4: GLUE results for models trained from scratch (FS) and from pretrained (CPT). Top row: From Scratch (FS), Bottom row: From Pretrained (CPT).
In the CPT setting (Figure 4, bottom row), we found that DEPTH and T5 perform similarly, both improving only slightly beyond the baseline. Initially, DEPTH seemed to perform worse than the baseline model, and did consistently worse than the T5-CPT model until later stages of pre-training. We speculate that this dip in performance is related to the change from the learned objective of the T5 model (from which DEPTH-CPT was initialized) from span-masking to span-masking and sentence un-shuffling. We share our full results on GLUE in Appendix E.
@@ 4.3 DiscoEval Fine-Tuning
We find that in the FS setting DEPTH consistently outperforms T5 across DC tasks, indicating its robustness in understanding discourse (Figure 5, top row). This suggests DEPTH’s pre-training objective is particularly beneficial for tasks that require a deep understanding of narrative structures (both in conversations as in DC-Chat, and more formal and informative texts as in DC-Wiki). We note that between steps 32k and 64k, DEPTH experienced a large positive boost in performance on DC-Wiki, perhaps indicative of an emergence [61] of discourse understanding during this phase of pre-training. For T5, we found that the model struggled to learn the SP-Arxiv task, achieving random-guess accuracy late in pre-training. However, in SP-Wiki and SP-Rocstory, T5 improves in performances between steps 64k and 128k, perhaps indicating an emergent ability occurring within this timeframe. We report our full results on DiscoEval in Appendix F.
@@
Figure 5: DiscoEval benchmark results for DEPTH and T5 models. Top row: From Scratch (FS), Bottom row: From Pretrained (CPT).
@@ 4.4 NI Fine-Tuning
In the FS setting (Figure 6(a)), we observe that DEPTH outperforms T5 in the NI benchmark, with a notable leap in performance between steps 16k and 32k. T5, by comparison, only improves significantly after step 64k, and obtains worse performance than DEPTH by the end of training. However, in the CPT setting (Figure 6(b)), DEPTH’s pre-training appears to hinder downstream performance compared to T5, possibly due to the domain shift from T5’s pre-training task to DEPTH’s pre-training task, which involves learning from shuffled inputs. We present a more complete analysis of these results in Appendix G.
@@ 5 Related Work
The potential of encoder-decoder architectures in today’s NLP landscape cannot be overstated. These architectures dominate context-heavy tasks ranging from translation [71, 65, 55] to summarization [67, 19, 55], and even following instructions across diverse domains [1, 60, 8]. Like their decoder-only counterparts, encoder-decoders are able to accommodate long inputs [19], and scale effectively effectively as a function of model size and training data [54]. Recently, Reka AI released a series of encoder-decoder models [41], where their dense 21B parameter model outperformed all models of its size in lmsys benchmark [70]  [^6] This Reka model is competitve with mixtral 8x22b [22] (which was trained with significantly more parameters using a mixture-of-experts architecture)..Encoder-deocder models are also strong multi-modal learners [41, 64, 15, 66]. When scaled sufficiently, encoder-decoders like Reca-Core may be competitive with state of the art models like GPT-4 [40], Gemini [56], and Claude-3.
While specialized encoder-decoder models such as PEGASUS [67], DialogVED [6], and a multi-party dialogue pre-training model [30] demonstrate the value of discourse-oriented tasks for encoder-decoder models, they have limited utility for broader tasks. Long-T5 [19] and UL2 [55] improved the ability of encoder-decoders to handle long contexts, but did not explicitly tackle discourse understanding. Flan-T5 [60] and Ex-T5 demonstrated the applicability of encoder-decoders across a variety of tasks, including ones that are heavily discourse dependent. However, these models depend on a vast yet costly annotated dataset to learn human preferences. The ERNIE series [68, 52, 53] integrates knowledge graphs and heuristic-guided masking for enriched representations but relies on extensive resources and faces challenges in timely knowledge updates. DEPTH, on the other hand, learns discourse through a generalized pre-training objective, and is not task-oriented or annotation-dependent.
@@ 6 Limitations
Given our lack of computational resources (Appendix C), we were not able to pre-train our models with a batch size that would allow an aggressive learning rate like that used in Raffel et al. [45]’s T5 (see Appendix A for additional details). We also pre-trained on substantially fewer tokens than T5. As a result, our model converges to a worse loss during pre-training, and performs worse on various downstream tasks. We also lacked sufficient computational resources to compute confidence intervals or statistical significance for our downstream experiments. [^7] Running a single downstream experiment on MNLI takes $5$-$7.5$ hours. We run $\approx 120$ experiments for each of 10 benchmarks, and do not have the capacity to repeat experiments 5$\geq$ 5$\times$ 5 × to obtain statistical significance.
Encoder-decoder LMs have fewer tools available for computationally efficient pre-training. For example, FlashAttention [11, 10], which provides a massive training speedup, is not available for encoder-decoder models. Furthermore, evolving approaches at the software level, such as Multi-Query Attention [47] for faster computation, and Ring-Attention [31] for extending context length, are not thoroughly vetted for encoder-decoder LMs. It is therefore difficult to create scalable pre-training experiments with new encoder-decoder architectures and objectives.
@@ 7 Conclusions and Future Work
Looking forward, the application of DEPTH to RAG [29], especially over sentence “chunks”, presents an exciting avenue for future research. Additionally, extending DEPTH’s pre-training objectives to encompass higher-level discourse units—such as paragraphs, chapters, and whole documents—offers further flexibility emerging hierarchical RAG systems [7]. Moreover, conducting further experiments with larger DEPTH models is helpful for understanding the scalability of discourse-focused training objectives. Such investigations could reveal whether the promising capabilities observed in DEPTH are amplified with increased model capacity.
@@ A.1 Our Speculations
We recognize that both DEPTH and T5 CPT models quickly reach a significantly lower reconstruction loss than their FS counterparts. While the original T5’s reached a final training loss of $\approx 0.75$ (with 512k steps of batch size 128 with packing) [45], our training loss reached $\approx 0.8$ (after 800k steps of batch size 200 without packing). We propose the following ideas to explain this gap:
@@ A.2 Example Packing
Given our choice of avoiding example-packing, we found that we were not able to pre-train T5 with the same hyper-parameters used in [45]. Specifically, we found that with a batch size of 128 and a learning rate of $1e^{-2}$, our model consistently diverged. This issue persisted with a learning rate of $1e^{-3}$. To stabilize our loss given the absence of packing, we used a lower maximum learning rate ($1e^{-4}$), which is in line with those used to pre-train BERT, SLM, and PMI. On the other hand, we see that given the same training parameters (i.e., learning rate and batch size), pre-training with packing can converge at a high learning rate (see Figure 7).
@@ Appendix C Computational Resources
We utilize 4 A40 GPUs, and 64 CPUs for training. We use a batch size of 200, since it helps us achieve much better GPU memory utilization. We leverage 16 CPUs for each GPU in order to increase the data loading time to accommodate for DEPTH’s more-complex corruption method, and to allow effective optimizer offloading with DeepSpeed Zero2 [46].
@@ D.2 Pre-Training Data
The rationale behind selecting C4 extends beyond its sheer volume and diversity. Given DEPTH’s architectural roots in the T5 model, which was originally pre-trained on C4, leveraging the same dataset facilitates a direct comparison of the enhancements our model introduces. This baseline compatibility is crucial for isolating the effects of our architectural and methodological innovations on the model’s performance. Furthermore, C4 is a subset of both Dolma [51] and RedPajama [9]. These datasets were used to train the most capable fully-open-source LMs (with released data) to date: OLMO [18] and LLama [57] respectively. This suggests that C4 is an effective component in pre-training effective LMs, and offers multiple additional datasets components that would be new to our model for future training. Finally, Muennighoff et al. [38] suggest that pre-training over the same data up to 4×4$\times$4 × still improves model performance, while [45]’s model has not completed even one full iteration over the C4 dataset. This hints at the viability of continuing to pre-train T5 on the same dataset it was already pre-trained on.
Appendix E GLUE Tables

### v2

@@ 3.1 Fine-tuning experiments
We follow up our pre-training experiments with a collection of downstream tasks. We evaluate our models on natural language inference (MNLI, Williams et al. (2018)), sentiment analysis (SST2, Socher et al. (2013)), and grammar (CoLA, Warstadt et al. (2019)) within the GLUE benchmark (Wang et al., 2018). We also use the DiscoEval suite (Chen et al., 2019) to evaluate models on their understanding of discourse. We use two tasks from DiscoEval: Sentence Permutation (SP) and Discourse Coherence (DC). SP involves identifying the correct position of a removed, while DC involves predicting whether or not a paragraph was coherent. Finally, we measure our model’s generative abilities on the Natural Instructions (NI) dataset (Mishra et al., 2022), which measures the ability of LMs to follow instructions, and served as a benchmark for NanoT5 (Nawrot, 2023).
Our experimental framework is inspired by Pythia (Biderman et al., 2023), which evaluates the performance of LMs on downstream tasks from intermediate checkpoints. We run evaluation with checkpoints from both T5 and DEPTH models, gathered at steps $\{2K,4K,\ldots,512K,1M\}$ in order to examine these models’ emergent capabilities. The exponential distance between these checkpoints allows us to scale intermediate checkpoint evaluation to much longer training runs. [^5] Evaluating intermediate checkpoint performance every 10,000 steps (as was done in Pythia) on datasets as large as MNLI is unfeasible with our limited computational resources.
@@ 4.1 C4 pre-training
During pre-training, we find that DEPTH consistently achieves a lower validation loss than a comparably trained T5 model. This is true for both FS and CPT. Furthermore, when we isolate the reconstruction loss (the objective used by T5, without sentence tokens), we find that DEPTH outperforms T5 despite balancing an additional pre-training objective (Figure 3 for FS and Figure 3 for CPT). These results are consistent with the findings in SLM, where their model converged faster, and on fewer tokens than models such as BERT and T5.
@@
While we were not able to match the performance of the baseline model of Raffel et al. (2020) (see Appendix A for speculations on why), we have obtained the lowest loss scores among PyTorch implementations of T5 models. Specifically, in our FS setting, we find that our randomly initialized T5 model outperforms the validation loss of NanoT5, achieving $1.65$ vs. $1.95$ at step 64,000.
@@ 4.2 GLUE fine-tuning
We found that over the course of FS pre-training, both models improved on GLUE tasks. However, T5’s improvement pattern was slower than DEPTH’s (top row of Figure 4). We found it difficult to replicate the results of the original T5 (both on GLUE tasks and the pre-training loss) as discussed in Appendix A. We project that with more substantial training (i.e., 1–3M pre-training steps, and $\geq 2048$ examples per batch, as in T5), DEPTH could match or exceed the performance of T5 and SLM on downstream tasks.
@@
Figure 4: GLUE results for FS and CPT models. Top row: From Scratch (FS), Bottom row: From Pretrained (CPT).
In the CPT setting (Figure 4, bottom row), we found that DEPTH and T5 perform similarly, both improving only slightly beyond the baseline. Fine-tuning DEPTH on early CPT checkpoints performs worse than fine-tuning comparable T5-CPT checkpoints. We speculate that this dip in performance is related to the change in objective from span-masking to span-masking and sentence un-shuffling. We share our full results on GLUE in Appendix F.
@@ 4.3 DiscoEval fine-tuning
We find that in the FS setting DEPTH consistently outperforms T5 across DC tasks, indicating its robustness in understanding discourse  [^6] In particular, sentence-level discourse relations, as discussed in Jernite et al. (2017) (Figure 5, top row). This suggests DEPTH’s pre-training objective is particularly beneficial for tasks that require a deep understanding of narrative structures (both in conversations as in DC-Chat, and more formal and informative texts as in DC-Wiki). We note that between steps 32k and 64k, DEPTH experienced a large positive boost in performance on DC-Wiki, perhaps indicative of an emergence (Wei et al., 2022b) of discourse understanding during this phase of pre-training. For T5, we found that the model struggled to learn the SP-Arxiv task, achieving random-guess accuracy late in pre-training. However, in SP-Wiki and SP-Rocstory, T5 improves in performances between steps 64k and 128k, perhaps indicating an emergent ability occurring within this timeframe. We report our full results on DiscoEval in Appendix G.
@@
Figure 5: DiscoEval results for DEPTH and T5 models. Top row: From Scratch (FS), Bottom row: From Pretrained (CPT).
@@ 4.4 NI fine-tuning
In the FS setting (Figure 6(a)), we observe that DEPTH outperforms T5 in the NI benchmark, with a notable leap in performance between steps 16k and 32k. T5, by comparison, only improves significantly after step 64k, and obtains worse performance than DEPTH by the end of training. However, in the CPT setting (Figure 6(b)), DEPTH’s pre-training appears to hinder downstream performance compared to T5, possibly due to the domain shift from T5’s pre-training task to DEPTH’s pre-training task, which involves learning from shuffled inputs. We present a more complete analysis of these results in Appendix H.
@@
4.5 Error Analysis
We performed error analysis on the DiscoEval benchmark to better understand the nature of discourse errors that DEPTH and T5 made. For the SP task, we show in Table 3 that DEPTH made more reasonable mistakes than T5. For example, in SP-Arxiv FS, 23% of DEPTH’s mistakes were reasonable, relative to 7% by T5). We define “reasonable” mistakes as incorrect predictions that would have still resulted in a coherent sentence ordering. Both models struggled with pronoun resolution, which frequently led to incorrect predictions (accounting for 10-30% of all predictions we observed). T5-FS, in particular, often failed to recognize when a removed sentence should come first, a mistake largely resolved in T5-CPT. We note that some examples correctly predicted by FS models were incorrectly predicted by CPT models, and vice versa.
In the DC task, we noted a significant number of incorrectly formatted predictions (e.g., “cooherent” rather than “coherent”), especially in the DC-Chat subset. Each of these incorrectly formatted predictions, when adjusted to a correctly formatted prediction, were incorrect (e.g., an example that the model predicted “cooherent” is labeled “incoherent”). We show in Table 4, that DEPTH-FS was incorrect in DC-Chat examples that humans might find ambiguous (i.e., replacing a random sentence leaves the resulting passage coherent), reinforcing its strength in handling more complex discourse structures. We discuss this further in Appendix E.
@@ 5 Related work
The potential of encoder-decoder architectures in today’s NLP landscape cannot be overstated. These architectures dominate context-heavy tasks ranging from translation Üstün et al. (2024); Xue et al. (2021); Tay et al. (2022) to summarization Zhang et al. (2020); Guo et al. (2022); Tay et al. (2022), and even following instructions across diverse domains Aribandi et al. (2022); Wei et al. (2022a); Chung et al. (2024). Like their decoder-only counterparts, encoder-decoders are able to accommodate long inputs (Guo et al., 2022), and scale effectively effectively as a function of model size and training data (Sutawika et al., 2024). Ormazabal et al. (2024) released a series of encoder-decoder models, where their dense 21B parameter model outperformed all models of its size in the lmsys benchmark (Zheng et al., 2023). [^7] This Reka model is competitve with mixtral 8x22b (Jiang et al., 2024) (which was trained with significantly more parameters using a mixture-of-experts architecture). Encoder-deocder models are also strong multi-modal learners (Ormazabal et al., 2024; Wu et al., 2023; Dosovitskiy et al., 2021; Zhai et al., 2022). When scaled sufficiently, encoder-decoders like Reca-Core may be competitive with state of the art models like GPT-4 (OpenAI et al., 2023), Gemini Team et al. (2023), and Claude-3.
While specialized encoder-decoder models such as PEGASUS (Zhang et al., 2020), DialogVED (Chen et al., 2022), and a multi-party dialogue pre-training model (Li et al., 2023) demonstrate the value of discourse-oriented tasks for encoder-decoder models, they have limited utility for broader tasks. Long-T5 (Guo et al., 2022) and UL2 (Tay et al., 2022) improved the ability of encoder-decoders to handle long contexts, but did not explicitly tackle discourse understanding. Flan-T5 (Wei et al., 2022a) and Ex-T5 (Aribandi et al., 2022) demonstrated the applicability of encoder-decoders across a variety of tasks, including ones that are heavily discourse dependent. However, these models depend on a vast yet costly annotated dataset to learn human preferences. Finally, BART (Lewis et al., 2020a) is an encoder-decoder which leverages sentence shuffling during pre-training, but does not train dedicated hierarchical representations for sentences (essentially behaving like a DEPTH model without sentence-tokens, and without attention-mask induced hierarchy).
@@ 6 Limitations
Given our lack of computational resources (Appendix C), we were not able to pre-train our models with a batch size that would allow an aggressive learning rate like that used in Raffel et al. (2020)’s T5 (see Appendix A for additional details). We also pre-trained on substantially fewer tokens than T5. As a result, our model converges to a worse loss during pre-training, and performs worse on downstream tasks. We also lacked computational resources to compute confidence intervals or statistical significance for our downstream experiments. [^8] Running a single downstream experiment on MNLI takes $5$-$7.5$ hours. We run $\approx 120$ experiments for each of 10 benchmarks, and do not have the capacity to repeat experiments 5$\geq$ 5$\times$ 5 × to obtain statistical significance.
Encoder-decoder LMs have fewer tools available for computationally efficient pre-training. For example, FlashAttention (Dao et al., 2022; Dao, 2023), which provides a massive training speedup, is not available for encoder-decoder models. It is therefore difficult to create scalable pre-training experiments with new encoder-decoder architectures and objectives.
@@ 7 Conclusions and future work
Looking forward, the application of DEPTH to RAG (Lewis et al., 2020b), especially over sentence “chunks”, presents an exciting avenue for future research. Additionally, extending DEPTH’s pre-training objectives to encompass higher-level discourse units—such as paragraphs, chapters, and whole documents—offers further flexibility emerging hierarchical RAG systems (Chen et al., 2024). Moreover, conducting further experiments with larger DEPTH models is helpful for understanding the scalability of discourse-focused training objectives. Such investigations could reveal whether the promising capabilities observed in DEPTH are amplified with increased model capacity. Finally, sentence-level pre-training tasks such as next-sentence prediction (as in Krishna et al. (2022) and Zhang et al. (2020)) may prove powerful alternatives to sentence un-shuffling.
Acknowledgments
This research was supported by the Israel Science Foundation (grant No. 448/20), an Azrieli Foundation Early Career Faculty Fellowship, and an AI Alignment grant from Open Philanthropy.
@@ A.1 Our speculations
We recognize that both DEPTH and T5 CPT models quickly reach a significantly lower reconstruction loss than their FS counterparts. While the original T5’s reached a final training loss of $\approx 0.75$ (with 512k steps of batch size 128 with packing) Raffel et al. (2020), our training loss reached $\approx 0.8$ (after 800k steps of batch size 200 without packing). We propose the following ideas to explain this gap:
@@ A.2 Example packing
Given our choice of avoiding example-packing, we found that we were not able to pre-train T5 with the same hyper-parameters used in Raffel et al. (2020). Specifically, we found that with a batch size of 128 and a learning rate of $1e^{-2}$, our model consistently diverged. This issue persisted with a learning rate of $1e^{-3}$. To stabilize our loss given the absence of packing, we used a lower maximum learning rate ($1e^{-4}$), which is in line with those used to pre-train BERT, SLM, and PMI. On the other hand, we see that given the same training parameters (i.e., learning rate and batch size), pre-training with packing can converge at a high learning rate (see Figure 7).
@@ Appendix C Computational resources
We utilize 4 A40 GPUs, and 64 CPUs for training. We use a batch size of 200, since it helps us achieve much better GPU memory utilization. We leverage 16 CPUs for each GPU in order to increase the data loading time to accommodate for DEPTH’s more-complex corruption method, and to allow effective optimizer offloading with DeepSpeed Zero2 (Rajbhandari et al., 2020).
@@ D.2 Pre-training data
The rationale behind selecting C4 extends beyond its sheer volume and diversity. Given DEPTH’s architectural roots in the T5 model, which was originally pre-trained on C4, leveraging the same dataset facilitates a direct comparison of the enhancements our model introduces. This baseline compatibility is crucial for isolating the effects of our architectural and methodological innovations on the model’s performance. Furthermore, C4 is a subset of both Dolma Soldaini et al. (2024) and RedPajama Computer (2023). These datasets were used to train the most capable fully-open-source LMs (with released data) to date: OLMO Groeneveld et al. (2024) and LLama Touvron et al. (2023) respectively. This suggests that C4 is an effective component in pre-training effective LMs, and offers multiple additional datasets components that would be new to our model for future training. Finally, Muennighoff et al. (2023) suggest that pre-training over the same data up to 4×4$\times$4 × still improves model performance, while (Raffel et al., 2020)’s model has not completed even one full iteration over the C4 dataset. This hints at the viability of continuing to pre-train T5 on the same dataset it was already pre-trained on.
Appendix E Error Analysis
We select randomly select 30 examples from each task in DiscoEval, and manually inspect the nature of our models’ errors. For example, in the SP task, we were interested in observing if a model tended to misunderstand clues from pronouns or transitions. For both DC and SP, we were also counted the number of examples where a human might find the LLM’s answer reasonable (given the human-perceived coherence of the example).
E.1 Sentence Position
Consistent with the macro-level results from our DiscoEval fine-tuning experiments, DEPTH generally performs better in FS than T5. However, in the CPT setting, while DEPTH shows better reasoning in some cases, T5-CPT outperforms it, particularly on simpler tasks like SP-Wiki and SP-Rocstory, likely benefiting from more consistent pretraining. DEPTH’s greater number of reasonable mistakes indicates its strength in engaging with complex discourse structures, but pronoun resolution and transition errors remain areas for further improvement.
Given the 30 examples we’ve sampled, we categorize the types of errors our models tend to make. An error type is a reason by which a person might be able to infer the correct label. If the model predicts incorrectly given an “obvious” hint (e.g. introducing an entity that is referenced via a pronoun), then we categorize the error type based on that hint. In Table 3 we show the counts of error types that each model made on each subset of SP.
Dataset Model 0 1 2 3 4
SP-Wiki CPT T5 0.57 0.07 0.23 0.13 0.07
DEPTH 0.57 0.03 0.23 0.2 0.03
SP-Wiki FS T5 0.57 0.07 0.23 0.13 0.07
DEPTH 0.57 0.03 0.23 0.2 0.03
SP-Arxiv CPT T5 0.63 0.07 0.2 0.1 0
DEPTH 0.57 0.13 0.2 0.1 0
SP-Arxiv FS T5 0.37 0.23 0.3 0.07 0.03
DEPTH 0.43 0.1 0.3 0.23 0
SP-Rocstory CPT T5 0.77 0.07 0.1 0.07 0
DEPTH 0.57 0.13 0.2 0.1 0
SP-Rocstory FS T5 0.47 0.2 0.27 0.07 0
DEPTH 0.6 0.2 0.13 0.07 0
Table 3: Approximate ratio of prediction types for SP-Wiki, SP-Arxiv, and SP-Rocstory across T5 and DEPTH models in FS and CPT settings. There are 30 total examples. Each prediction over these examples is categorized into one of 5 prediction types: 0 - Correct prediction, 1 - Incorrect (hint from transitions), 2 - Incorrect (hint from pronoun), 3 - Incorrect (reasonable error), 4 - Incorrect (hint from punctuation).
E.2 Discourse Coherence
We found that in the Discourse Coherence (DC) subset, both models were strictly incorrect by predicting poorly formatted outputs. Each of the poorly formatted predictions was incorrect (e.g., if the model predicted “cooherent” instead of “coherent”, the correct label was “incoherent”). We analyze the ratio of these types of predictions on DC in Table 4. Further, we find that in DC-Chat, both models tend to make errors that a human might find reasonable. This implies that the augmentation on the input example (i.e., replacing one if the sentences in the paragraph with another one) did not adversely impact the example’s coherence. Strangely, we found that FS models predicted more correct outputs than their CPT counterparts. This may be the result of selecting a too small a sample-size of examples to analyze.
Dataset Model 0 1 2
DC-Wiki-FS T5 0.87 0.32 0.00
DEPTH 0.87 0.13 0.00
DC-Wiki-CPT T5 0.80 0.08 0.00
DEPTH 0.83 0.08 0.00
DC-Chat-FS T5 0.60 0.42 0.17
DEPTH 0.63 0.35 0.23
DC-Chat-CPT T5 0.83 0.30 0.10
DEPTH 0.77 0.29 0.07
Table 4: Prediction types for models and dataset splits. 0: Correct predictions (Type 0), 1: Poorly formatted predictions (Type 1), 2: Reasonable errors (Type 2).
Appendix F GLUE results